b=10a-60
a-6=10b is this correct or incorrect equation and tell why or why not
I am 15 and home schooling and I really need help

Found 2 solutions by jankiz, Alan3354:
Answer by jankiz(13)

It is incorrect, because b=10a-60 does not equal to a-6=10b. You added 10 to b, which you can't because they are not like terms. The correct equation if you are solving for b is b/10=a-6. If you are solving for a, then it would be b/10+6=a.

Answer by Alan3354(69443)

b=10a-60
--------------
If you want to solve for a:
10a - 60 = b
Add 60
10a = b + 60
Divide by 10
a = (b + 60)/10
or,
a = (b/10) + 6 same thing